If f(x)=x^2+3x, find f(2-a)
Answer by unlockmath(1688)

Hello,
If f(x)=x^2+3x, find f(2-a)
This is rewritten as:
f(x)=(2-a)^2+3(2-a):
Expand it out:
F(x)=4-4a+a^2+6-3a
Combine like terms:
F(x)=a^2-7a+10
Factor:
(a-5)(a-2)
a=5
a=2
Make sense?
RJ
www.math-unlock.com